OpenStudy (anonymous):

write an equation of a circle with the given center and radius center (-7 -6) and radius 2

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

if center (a,b) and radius is r then equation of the circle is \[(x-a)^2 +(y-b)^2= r^2\]

OpenStudy (anonymous):

so it would be (x+7)^2+(y+6)^2=4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome @Jelia97